# Learning with Errors over Group Rings Constructed by Semi-direct ${}^{\dagger}$

### 01 Dec 2023 v2

@@ I-C Applications
Informally, let $R$ be some group ring or a quotient ring of some group ring. For instance, we can select a ring $R=\mathbb{Z}[G]/\langle t^{n/2}+1\rangle$, where $G$ is a semi-direct product of cyclic groups $\mathbb{Z}_{m}=\langle s\rangle$ and $\mathbb{Z}_{n}=\langle t\rangle$. Let $q$ be a positive integer. The public-key cryptosystem based on $\mathsf{GR}\mbox{-}\mathsf{LWE}$ is as follows: The key-generation algorithm first generates a uniformly random element $a\in R_{q}:=R/qR$ (under coefficient embedding) and two “short” elements $s,e\in R_{q}$ from error distribution (typically selected from a discrete Gaussian distribution with “narrow” error). It outputs a $\mathsf{GR}\mbox{-}\mathsf{LWE}$ sample $(a,b=s\cdot a+e)\in R_{q}\times R_{q}$ as the public key and $s$ as the secret key. To encrypt a $|G|/2$-bit messages $m\in\{0,1\}^{|G|/2}$ that can be viewed as a polynomial of degree $n/2-1$ with 00 or $1$ coefficients, the encryption algorithm first generates three “short” random elements $r,e_{1}$ and $e_{2}\in R_{q}$ from error distribution and outputs $(u,v)\in R_{q}\times R_{q}$ with $u=a\cdot r+e_{1}\bmod{q}$ and $v=b\cdot r+e_{2}+\lfloor q/2\rceil\cdot m\bmod{q}.$ To decrypt $(u,v)\in R_{q}\times R_{q}$, the decrypt algorithm computes
@@
By choosing appropriate parameters, the element $e\cdot r-s\cdot e_{1}+e_{2}$ can be restricted to be very “short”. By rounding the coefficient of $\hat{m}$ to either 00 or $\lfloor q/2\rceil$, we can efficiently recover the bit string except with negligible probability.
@@ I-D Paper organization
In Section 2, we give some basic knowledge of lattice and related computaionally hard problems. In Section 3, we provide a reduction from worst-case lattice problems to the search version of $\mathsf{GR}\mbox{-}\mathsf{LWE}$ over some group rings with special properties. In Section 4, we present a direct reduction from worst-case lattice problems to the decision version of $\mathsf{GR}\mbox{-}\mathsf{LWE}$, considering two types of group rings, This generalizes the results of [30] and enables to construct some cryptographic primitives. In Section 5, we summarize our findings and offer some thoughts for further study.
@@ II Preliminaries
If $S$ is a nonempty set, then $x\leftarrow S$ denotes that $x$ is a random variable uniformly distributed in $S$. If $\varphi$ is a probability density function, then $x\leftarrow\varphi$ denotes that $x$ is a random variable distributed as $\varphi$. Given two probability distribution functions $\rho_{1},\rho_{2}$ over $\mathbb{R}^{n}$, the statistical distance between $\rho_{1}$ and $\rho_{2}$ is defined as
[…]
@@ II-A Lattice Background
[…]
@@
:=$\mathcal{L}^{\vee}:=\left\{\mathbf{x}\in\mathbb{R}^{n}:\langle\mathbf{x},\mathbf{y}\rangle\in\mathbb{Z},\text{ for all }\mathbf{y}\in\mathcal{L}\right\},$
[…]
@@ II-B Gaussian measures
Definition 1 (Smoothing parameter, [45, 30]).
@@
$\sum_{\mathbf{y}\in\mathcal{L}^{\vee}\backslash\{0\}}D_{1/s}(\mathbf{y})=\sum_{\mathbf{y}\in\mathcal{L}^{\vee}\backslash\{0\}}\exp(-\pi s^{2}\cdot\|\mathbf{y}\|^{2})\leq\varepsilon.$
@@
$\sum_{\mathbf{y}\in\mathcal{L}^{\vee}\backslash\{0\}}\exp(-\pi\cdot\mathbf{y}^{t}\mathbf{A}\mathbf{A}^{t}\mathbf{y})\leq\varepsilon,$
@@
For any $n$-dimensional lattice $\mathcal{L}$, we have $\eta_{2^{-2n}}(\mathcal{L})\leq\sqrt{n}/\lambda_{1}(\mathcal{L}^{\vee})$, and $\eta_{\varepsilon}(\mathcal{L})\leq\sqrt{\ln(n/\varepsilon)}\lambda_{n}(\mathcal{L})$ for all $10<\varepsilon<$ 1.
@@ II-D Group ring
Let $G=\{g_{1},g_{2},\cdots,g_{n}\}$ be a finite group of order $n$. The elements in group ring $R[G]$ are formal sums
@@
Addition is defined as
@@
Multiplication is defined as
$\left(\sum\limits_{i=1}^{n}a_{i}g_{i}\right)\left(\sum\limits_{i=1}^{n}b_{i}g_{i}\right)=\sum\limits_{r=1}^{n}\left(\sum\limits_{g_{i}g_{j}=g_{r}}a_{i}b_{j}\right)g_{r}.$
Naturally, $R[G]$ can be seen as a free $R$-module of rank $|G|=n$, and $\{g_{1},g_{2},\ldots,g_{n}\}$ naturally forms a $R$-basis.
@@ II-E Semi-direct product
Let $G$ and $H$ be two groups, and $\varphi:G\rightarrow\operatorname{Aut}(H)$ be a group homomorphism, where $\operatorname{Aut}(H)$ represents the automorphism group of $H$. The semi-direct product group of $G$ and $H$ induced by $\varphi$ is defined as a set $K=\{(g,h)\mid g\in G,h\in H\}$ with multiplication law “normal-$\cdot$”:
@@
We denote $K$ with multiplication “normal-$\cdot$” by $G\ltimes_{\varphi}H$, or simply $G\ltimes H$.
@@
Then $D_{2n}$ can also be represent as $\mathbb{Z}_{2}\ltimes_{\varphi}\mathbb{Z}_{n}$, where $\varphi(s)=\operatorname{inv}$, where $\operatorname{id}$ maps each element of $\mathbb{Z}_{n}$ to its inverse (under addition modulo $n$).
@@
Similar to Type I, the semi-direct product of $\mathbb{Z}_{n}^{\ast}$ and $\mathbb{Z}_{n}$ induced by $\psi$ can be expressed as follows:
@@
    and $r(u-1)$ (non-equivalent) 2-dimensional irreducible representations $(i=1,2,\ldots,r$, $j=1,2,\ldots,u-1)$ satisfying
@@ II-F Ideal lattice and coefficient embedding
[…]
@@
Through the study of ideal lattice $\mathcal{I}$ of a group ring, it is crucial to consider the relationship between dual ideal lattice $\mathcal{I}^{\vee}$ and inverse ideal $\mathcal{I}^{-1}$. In the following lemma, we primarily consider the two types of groups defined by (2) and (3). We show that the dual ideal lattice and the inverse ideal lattice of the same invertible ideal lattice are equivalent, up to a specific permutation of the coordinates. The following lemma generalized the result of Lemma 3 of [30] to Type I and Type II. The proof is attached in the appendix.
@@ II-G Lattice problems
We introduce some important and useful problems which are believed to be computionally hard. They are commonly used to characterize the hardness of $\mathsf{LWE}$ variants. These problems include Shortest Vector Problem ($\mathsf{SVP}$), Shortest Independent Vectors Problem ($\mathsf{SIVP}$), Closest Vector Problem ($\mathsf{CVP}$), and Bounded Distance Decoding ($\mathsf{BDD}$).
@@ II-H Natural inclusion mapping
Let $q$ be a positive integer and $\mathcal{L}$ be a lattice. Denote $\mathcal{L}_{q}$ as the quotient $\mathcal{L}/q\mathcal{L}$. For any lattices $\mathcal{L}^{\prime}\subseteq\mathcal{L}$, the natural inclusion map is defined as $\varphi:\mathcal{L}_{q}^{\prime}\rightarrow\mathcal{L}_{q}$ which maps $x+q\mathcal{L}^{\prime}$ to $x+q\mathcal{L}$ [^3] It can be verified that the definition of natural inclusion map is well-defined under the condition that $\mathcal{L}^{\prime}\subseteq\mathcal{L}$. The natural inclusion map can be viewed as a composition of a natural homomorphism $\mathcal{L}/q\mathcal{L}^{\prime}\rightarrow(\mathcal{L}/q\mathcal{L}^{\prime})/(q\mathcal{L}/q\mathcal{L}^{\prime})=\mathcal{L}/q\mathcal{L}$ and an inclusion map $\mathcal{L}^{\prime}/q\mathcal{L}^{\prime}\rightarrow\mathcal{L}/q\mathcal{L}^{\prime}$. The following lemma is an important result from [13]. It describes under what condition the natural inclusion map $\varphi$ is a bijection. We include the proof for completeness.
@@
Proof.
Let $\mathbf{b}\in\mathcal{L}^{n},\mathbf{b}^{\prime}\in\mathcal{L}^{\prime n}$ be some bases of $\mathcal{L}$ and $\mathcal{L}^{\prime}$ as free $\mathbb{Z}$-modules, respectively. It is straightforward to see that $\mathbf{b},\mathbf{b}^{\prime}$ are also bases of $\mathcal{L}_{q}$ and $\mathcal{L}^{\prime}_{q}$ as free $\mathbb{Z}_{q}$-modules. Since $\mathcal{L}^{\prime}$ is a subset of $\mathcal{L}$, then there exists a square matrix $\mathbf{T}\in\mathbb{Z}^{n\times n}$, such that $\mathbf{b}^{\prime}=\mathbf{T}\cdot\mathbf{b}.$ Let $\mathbf{x}^{\prime}\in\mathbb{Z}_{q}^{n}$ be the coefficient vector of some $x^{\prime}\in\mathcal{L}_{q}^{\prime}$ corresponding to the $\mathbb{Z}_{q}$-basis $\mathbf{b}^{\prime}$. We have
$x^{\prime}=\langle\mathbf{b}^{\prime},\mathbf{x}^{\prime}\rangle=\langle\mathbf{T}\cdot\mathbf{b},\mathbf{x}^{\prime}\rangle=\langle\mathbf{b},\mathbf{T}^{t}\cdot\mathbf{x}^{\prime}\rangle.$
Thus, $\mathbf{x}=\mathbf{T}^{t}\cdot\mathbf{x}^{\prime}\in\mathbb{Z}_{q}^{n}$ is the coefficient vector of the image of $x^{\prime}$, i.e., $\varphi(x^{\prime})\in\mathcal{L}_{q}$ related to $\mathbb{Z}_{q}$-basis $\mathbf{b}$. From the above, we can conclude that the natural inclusion map $\varphi$ is exactly determined by the left multiplication of $\mathbf{T}$. Then $\varphi$ is bijective if and only if $\mathbf{T}$ is invertible over $\mathbb{Z}_{q}$ if and only if $q$ is coprime with $|\det(\mathbf{T})|=|\mathcal{L}/\mathcal{L}^{\prime}|.$ Furthermore, by calculating $\mathbf{T}$, we can operate $\varphi$ and $\varphi^{-1}$ efficiently. ∎
@@
We can also apply Lemma 11 to dual lattice $\mathcal{L}^{\vee}$. As $|\mathcal{L}/\mathcal{L}^{\prime}|=|(\mathcal{L}^{\prime})^{\vee}/\mathcal{L}^{\vee}|$, we can obtain that
@@ II-I Group ring $\mathsf{LWE}$
Here we describe the $\mathsf{LWE}$ problem over group rings. We use the following notations. Let $G$ be a finite group and let $R$ be the group ring $\mathbb{Z}[G]$ itself or one of its quotient rings. For an integer modulus $q\geq 2$, let $R_{q}$ denote the quotient ring $R/qR$. Likewise, for any (fractional) ideal $\mathcal{I}$ of a ring $R$, let $\mathcal{I}_{q}$ denote $\mathcal{I}/q\mathcal{I}$. We also denote $R_{\mathbb{R}}$ as the tensor product $R\otimes_{\mathbb{Z}}\mathbb{R}$ and let $\mathbb{T}:=R_{\mathbb{R}}/R$.
@@
Definition 9 (Search Group ring-$\mathsf{LWE}$).
Let $\Psi$ be a family of distributions over $R_{\mathbb{R}}\times\mathbb{T}$. The search version Group ring-$\mathsf{LWE}$ problem asks to find the secret element $s\in R_{q}$, given arbitrarily many independent samples $(a_{i},b_{i})\in R_{q}\times\mathbb{T}$ chosen from the $R\mbox{-}\mathsf{LWE}$ distribution $A_{s,\psi}$, where $\psi$ is a distribution of $\Psi$. We denote such search problem by $R\mbox{-}\mathsf{LWE}_{q,\Psi}$.
Definition 10 (Average-case decision Group ring-$\mathsf{LWE}$).
@@ III-C The $\mathsf{GDP}$ to Search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ reduction
$\sum\limits_{\mathbf{y}\in\mathcal{L}^{\vee}\backslash\{0\}}\exp(-\pi\mathbf{y}^{t}(\mathbf{A}^{-t}\mathbf{A}^{-1}+\dfrac{1}{s^{2}}\mathbf{B}^{t}\mathbf{B})^{-1}\mathbf{y})\leq\varepsilon$
@@
By applying this lemma to a fractional ideal in $\mathbb{R}[G]$ (where the elements are under coefficient embedding), we obtain the following corollary.
@@
Let $R=\mathbb{Z}[G]$ be a group ring, where $G$ is a group of Type I or Type II with $n$ elements. Let $\alpha>0$ be a real, let $q>2$ be an integer, and let $r>\sqrt{2}q\cdot\eta_{\varepsilon}(\mathcal{I})$ be a real. Then given an invertible ideal $\mathcal{I}$ of $R$ with index $|R/\mathcal{I}|=\det(\mathcal{I})$ coprime with $q$, there exists an efficient algorithm that given an instance from $\mathsf{BDD}_{\mathcal{I}^{-1},d}$ (in matrix norm), where $\mathcal{I}^{-1}$ is the left inverse of $\mathcal{I}$ and samples from $D_{\mathcal{I},r}$, outputs samples that have $R$-$\mathsf{LWE}$ distribution $A_{q,\psi}$ (up to negligible statistical distance) for some $\psi\in\Psi_{\leq\alpha}$, where $d=\alpha q/(\sqrt{2}r)$.
@@
For an $\mathsf{BDD}_{\mathcal{I}^{-1},d}$ instance $y=x+e$ where $x$ is an element of $\mathcal{I}^{-1}$ and the matrix norm of $e$ is bounded by $d$. We can construct an $R\mbox{-}\mathsf{LWE}$ sample as follows.
@@
$b=(y\cdot z)/q+e^{\prime}\bmod R=(x\cdot z)/q+(e\cdot z)/q+e^{\prime}\bmod R\in 1/q\cdot(R/qR),$

### 10 Mar 2026 v3

@@ I-C Applications
Informally, let $R$ be some group ring or a quotient ring of some group ring. For instance, we can select a ring $R=\mathbb{Z}[G]/\langle t^{n/2}+1\rangle$, where $G$ is a semi-direct product of cyclic groups $\mathbb{Z}_{m}=\langle s\rangle$ and $\mathbb{Z}_{n}=\langle t\rangle$. Let $q$ be a positive integer. The public-key cryptosystem based on $\mathsf{GR}\mbox{-}\mathsf{LWE}$ is as follows: The key-generation algorithm first generates a uniformly random element $a\in R_{q}:=R/qR$ (under coefficient embedding) and two “short” elements $s,e\in R_{q}$ from error distribution (typically selected from a discrete Gaussian distribution with “narrow” error). It outputs a $\mathsf{GR}\mbox{-}\mathsf{LWE}$ sample $(a,b=s\cdot a+e)\in R_{q}\times R_{q}$ as the public key and $s$ as the secret key. To encrypt a $|G|/2$-bit message $m\in\{0,1\}^{|G|/2}$ that can be viewed as a polynomial of degree n/21n/2-1 with 0 or $1$ coefficients, the encryption algorithm first generates three “short” random elements $r,e_{1}$ and $e_{2}\in R_{q}$ from error distribution and outputs $(u,v)\in R_{q}\times R_{q}$ with $u=a\cdot r+e_{1}\bmod{q}$ and $v=b\cdot r+e_{2}+\lfloor q/2\rceil\cdot m\bmod{q}.$ To decrypt $(u,v)\in R_{q}\times R_{q}$, the decrypt algorithm computes
@@
By choosing appropriate parameters, the element $e\cdot r-s\cdot e_{1}+e_{2}$ can be restricted to be very “short”. By rounding the coefficient of $\hat{m}$ to either 0 or $\lfloor q/2\rceil$, we can efficiently recover the bit string except with negligible probability.
@@ I-D Paper organization
In Section 2, we give some basic knowledge of lattice and related computationally hard problems. In Section 3, we provide a reduction from worst-case lattice problems to the search version of $\mathsf{GR}\mbox{-}\mathsf{LWE}$ over some group rings with special properties. In Section 4, we present a direct reduction from worst-case lattice problems to the decision version of $\mathsf{GR}\mbox{-}\mathsf{LWE}$, considering two types of group rings. This generalizes the results of [30] and enables to construct some cryptographic primitives. In Section 5, we summarize our findings and offer some thoughts for further study.
@@ II Preliminaries
If $S$ is a nonempty finite set, then $x\leftarrow S$ denotes that $x$ is a random variable uniformly distributed over $S$. If $\varphi$ is a probability density function, then $x\leftarrow\varphi$ denotes that $x$ is a random variable distributed as $\varphi$. Given two probability distribution functions $\rho_{1},\rho_{2}$ over $\mathbb{R}^{n}$, the statistical distance between $\rho_{1}$ and $\rho_{2}$ is defined as
$\Delta(\rho_{1},\rho_{2}):=\frac{1}{2}\int_{\mathbb{R}^{n}}|\rho_{1}(\mathbf{x})-\rho_{2}(\mathbf{x})|d\mathbf{x}.$
@@ II-A Lattice Background
The minimum distance of a lattice $\mathcal{L}$ in a given norm $\|\cdot\|$ is the length of a shortest nonzero lattice vector, i.e.,
@@
$\mathcal{L}^{\vee}:=\left\{\mathbf{x}\in\mathbb{R}^{n}:\langle\mathbf{x},\mathbf{y}\rangle\in\mathbb{Z},\text{ for all }\mathbf{y}\in\mathcal{L}\right\},$
where $\langle\cdot,\cdot\rangle$ represents inner product (Euclidean inner product by default). And it is easy to prove that $\mathcal{L}^{\vee}$ is also a lattice in $\mathbb{R}^{n}$ and if $\mathbf{B}$ is a basis of $\mathcal{L}$, then $\mathbf{B}^{-t}:=(\mathbf{B}^{-1})^{t}=(\mathbf{B}^{t})^{-1}$ is a basis of $\mathcal{L}^{\vee}$, hence $(\mathcal{L}^{\vee})^{\vee}=\mathcal{L}.$
@@ II-B Gaussian measures
Definition 1 (Smoothing parameter,[30], [46]).
@@
[…]
@@
[…]
@@
For any $n$-dimensional lattice $\mathcal{L}$, we have $\eta_{2^{-2n}}(\mathcal{L})\leq\sqrt{n}/\lambda_{1}(\mathcal{L}^{\vee})$, and $\eta_{\varepsilon}(\mathcal{L})\leq\sqrt{\ln(n/\varepsilon)}\lambda_{n}(\mathcal{L})$ for all $0<\varepsilon<1$.
@@ II-D Group ring
We follow the formulation of this definition as presented in [30]. Let $G=\{g_{1},g_{2},\cdots,g_{n}\}$ be a finite group of order $n$, and let $R$ be a commutative ring with identity. The group ring $R[G]$ is the set of all finite formal sums
@@
Addition is defined coefficient-wise
@@
Multiplication is defined by extending the group operations $R$-bilinearly:
$\left(\sum\limits_{i=1}^{n}a_{i}g_{i}\right)\left(\sum\limits_{i=1}^{n}b_{i}g_{i}\right)=\sum\limits_{r=1}^{n}\left(\sum\limits_{(i,j):\ g_{i}g_{j}=g_{r}}a_{i}b_{j}\right)g_{r}.$
In particular, $R[G]$ is a free $R$-module of rank $|G|=n$ in a natural way, with $\{g_{1},g_{2},\ldots,g_{n}\}$ as an $R$-basis.
@@ II-E Semi-direct product
Let $G$ and $H$ be two groups, and $\varphi:G\rightarrow\operatorname{Aut}(H)$ be a group homomorphism, where $\operatorname{Aut}(H)$ represents the automorphism group of $H$. The semi-direct product group of $G$ and $H$ induced by $\varphi$ is defined as a set $K=\{(g,h)\mid g\in G,h\in H\}$ with multiplication law $\cdot$”:
@@
We denote $K$ with multiplication $\cdot$” by $G\ltimes_{\varphi}H$, or simply $G\ltimes H$.
@@
Then $D_{2n}$ can also be represent as $\mathbb{Z}_{2}\ltimes_{\varphi}\mathbb{Z}_{n}$, where $\varphi(s)=\operatorname{inv}$, where $\operatorname{inv}$ maps each element of $\mathbb{Z}_{n}$ to its inverse (under addition modulo $n$).
@@
Similar to Type I, the semi-direct product of $\mathbb{Z}_{n}^{\ast}$ and $\mathbb{Z}_{n}$ can be induced by $\psi$. Here we use the symbol $a\odot g^{k}$ to represent the element $(a,g^{k})$ in the group $\mathbb{Z}_{n}^{\ast}\ltimes\mathbb{Z}_{n}$. Then the semi-direct product can be expressed as follows:
@@
    and r(u1)r(u-1) (non-equivalent) 2-dimensional irreducible representations $(i=1,2,\ldots,r$, $j=1,2,\ldots,u-1)$ satisfying
@@ II-F Ideal lattice and coefficient embedding
where $\|\cdot\|$ on the right-hand side is any norm defined on vectors in $\mathbb{R}^{n}$.
@@
Through the study of ideal lattice $\mathcal{I}$ of a group ring, it is crucial to consider the relationship between dual ideal lattice $\mathcal{I}^{\vee}$ and inverse ideal $\mathcal{I}^{-1}$. In the following lemma, we primarily consider the two types of groups defined by (2) and (3). We show that the dual ideal lattice and the inverse ideal lattice of the same invertible ideal lattice are equivalent, up to a specific permutation of the coordinates. The following lemma generalizes the result of Lemma 3 of [30] to Type I and Type II. The proof is attached in the appendix.
@@ II-G Lattice problems
We introduce some important and useful problems which are believed to be computionally hard. They are commonly used to characterize the hardness of $\mathsf{LWE}$ variants. These problems include Shortest Vector Problem ($\mathsf{SVP}$), Shortest Independent Vectors Problem ($\mathsf{SIVP}$), Closest Vector Problem ($\mathsf{CVP}$), and Bounded Distance Decoding ($\mathsf{BDD}$). The following definitions follow the standard formulations used in [8].
@@ II-H Natural inclusion mapping
Let $q$ be a positive integer and $\mathcal{L}$ be a lattice. Denote $\mathcal{L}_{q}$ as the quotient $\mathcal{L}/q\mathcal{L}$. For any lattices $\mathcal{L}^{\prime}\subseteq\mathcal{L}$, the natural inclusion map is defined as $\varphi:\mathcal{L}_{q}^{\prime}\rightarrow\mathcal{L}_{q}$ which maps $x+q\mathcal{L}^{\prime}$ to $x+q\mathcal{L}$. [^3] It can be verified that the definition of natural inclusion map is well-defined under the condition that $\mathcal{L}^{\prime}\subseteq\mathcal{L}$. The natural inclusion map can be viewed as a composition of a natural homomorphism $\mathcal{L}/q\mathcal{L}^{\prime}\rightarrow(\mathcal{L}/q\mathcal{L}^{\prime})/(q\mathcal{L}/q\mathcal{L}^{\prime})=\mathcal{L}/q\mathcal{L}$ and an inclusion map $\mathcal{L}^{\prime}/q\mathcal{L}^{\prime}\rightarrow\mathcal{L}/q\mathcal{L}^{\prime}$.The following lemma is an important result from [13]. It describes under what condition the natural inclusion map $\varphi$ is a bijection.
@@
[…]
@@ II-I Group ring $\mathsf{LWE}$
Here we describe the $\mathsf{LWE}$ problem over group rings. We use the following conventional notations from [1, 30]. Let $G$ be a finite group and let $R$ be the group ring $\mathbb{Z}[G]$ itself or one of its quotient rings. For an integer modulus $q\geq 2$, let $R_{q}$ denote the quotient ring $R/qR$. Likewise, for any (fractional) ideal $\mathcal{I}$ of a ring $R$, let $\mathcal{I}_{q}$ denote $\mathcal{I}/q\mathcal{I}$. We also denote $R_{\mathbb{R}}$ as the tensor product $R\otimes_{\mathbb{Z}}\mathbb{R}$ and let $\mathbb{T}:=R_{\mathbb{R}}/R$.
@@
Definition 9 ([30], Search Group ring-$\mathsf{LWE}$).
Let $\Psi$ be a family of distributions over $R_{\mathbb{R}}\times\mathbb{T}$. The search version Group ring-$\mathsf{LWE}$ problem asks to find the secret element $s\in R_{q}$, given arbitrarily many independent samples $(a_{i},b_{i})\in R_{q}\times\mathbb{T}$ chosen from the $R\mbox{-}\mathsf{LWE}$ distribution $A_{s,\psi}$ for some arbitrary $s\in R_{q}$, where $\psi$ is a distribution of $\Psi$. We denote such search problem by $R\mbox{-}\mathsf{LWE}_{q,\Psi}$.
Definition 10 ([30], Average-case decision Group ring-$\mathsf{LWE}$).
@@ III-C The $\mathsf{GDP}$ to Search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ reduction
[…]
@@
By applying this lemma to a fractional ideal in $\mathbb{R}[G]$ (where the elements are under coefficient embedding), we obtain the following corollary, which generalizes Corollary 1 of [30].
@@
Let $R=\mathbb{Z}[G]$ be a group ring, where $G$ is a group of Type I or Type II with $n$ elements. Let $\alpha>0$ be a real, let $q>2$ be an integer, and let $r>\sqrt{2}q\cdot\eta_{\varepsilon}(\mathcal{I})$ be a real. Then given an invertible right ideal $\mathcal{I}$ of $R$ with index $|R/\mathcal{I}|=\det(\mathcal{I})$ coprime with $q$, there exists an efficient algorithm that given an instance from $\mathsf{BDD}_{\mathcal{I}^{-1},d}$ (in matrix norm), where $\mathcal{I}^{-1}$ is the left inverse of $\mathcal{I}$ and samples from $D_{\mathcal{I},r}$, outputs samples that have $R$-$\mathsf{LWE}$ distribution $A_{q,\psi}$ (up to negligible statistical distance) for some $\psi\in\Psi_{\leq\alpha}$, where $d=\alpha q/(\sqrt{2}r)$.
@@
For an $\mathsf{BDD}_{\mathcal{I}^{-1},d}$ instance $y=x+e$ where $x$ is an element of $\mathcal{I}^{-1}$ and the matrix norm of $e$ is bounded by $d$, we can construct an $R\mbox{-}\mathsf{LWE}$ sample as follows.
@@
$b=(y\cdot z)/q+e^{\prime}\bmod R=(x\cdot z)/q+(e\cdot z)/q+e^{\prime}\bmod R,$